# Designing Adaptive User Interfaces for mHealth applications targeting chronic disease: A User-Centric Approach

### v1

@@ Abstract.
mHealth interventions show significant potential to help in the self-management of chronic diseases, but their under use remains a problem. Considering the substantial diversity among individuals dealing with chronic diseases, tailored strategies are essential. Adaptive User Interfaces (AUIs) may help address the diverse and evolving needs of this demographic. To investigate this approach, we developed an AUI prototype informed by existing literature findings. We then used this prototype as the basis for focus group discussions and interview studies with 22 participants managing various chronic diseases, and follow-up surveys of all participants. Through these investigations, we pinpointed key challenges related to the use of AUIs, strategies to improve adaptation design, and potential trade-offs between these challenges and strategies. Concurrently, a quantitative survey was conducted to extract preferences for AUIs in chronic disease-related applications with 90 further participants. This uncovered participants’ preferences for various adaptations, data types, collection methods, and involvement levels. Finally, we synthesised these insights and categories, aligning them with existing guidelines and design considerations for mHealth app adaptation design. This resulted in nine guidelines that we refined by a final feedback survey conducted with 20 participants.
@@ 1. Introduction
Chronic diseases, including conditions such as asthma, heart disease, and diabetes, present formidable challenges to health care systems around the world (WHO, 2022). The management of these long-term health conditions transcends simple medical treatment, with an increasing emphasis on empowering patients to actively engage in self-management practices (WHO, 2022). The use of mHealth technologies has emerged as a promising avenue for promoting self-management by strengthening medication adherence and facilitating self-tracking capabilities (Hamine et al., 2015). However, despite the potential of these technologies to improve health outcomes, research indicates that a significant proportion of people who benefit from mHealth solutions do not fully take advantage of them (Han et al., 2010). In order to increase the adoption of mHealth applications, it is essential to develop more user-friendly systems capable of meeting the varied needs of users (Han et al., 2010). However, several challenges need to be addressed to achieve this objective. Firstly, chronic diseases are inherently heterogeneous, impacting individuals in various ways, including triggers, symptoms, and severity (Harvey et al., 2012; Audulv, 2013). Second, it is crucial to consider the development stages of chronic diseases when designing mHealth technology, as chronic diseases can cooccur with other medical or psychological disorders, further adding to the complexity of self-management of chronic diseases (di San Marco et al., 2019; Audulv, 2013). Third, chronic diseases generally persist over an individual’s lifetime (WHO, 2022; Harvey et al., 2012). Therefore, mHealth technologies are needed to keep users engaged and motivated in the long run. This makes chronic disease management beyond the narrow definition of disease management.
Adaptive User Interfaces (AUIs) have often been recommended to accommodate the considerable variability between patients by tailoring the user interface (UI) to align with individual needs, objectives and usage contexts in mHealth technologies (Norcio and Stanley, 1989). Despite increasing interest in employing AUIs within chronic disease related applications (Setiawan et al., 2019; Talboom-Kamp et al., 2018), they often overlook diverse user characteristics and interactions (Grua et al., 2020; McLean et al., 2011). Furthermore, there remains a dearth of information regarding the development processes for AUIs during the early design phases (Wang et al., 2023b). Current AUI evaluation methods focus mainly on evaluating the overall effectiveness of the application, making it difficult to determine the utility of AUIs and user preferences (Wang et al., 2023b). Furthermore, there is limited information on how people with chronic diseases use AUIs, potentially resulting in the underutilisation of adaptation benefits (Kornfield et al., 2020). Addressing these gaps is crucial to optimising AUI design and improving user engagement in chronic disease management applications.
In this work, we examine the perspectives of users on the usage of AUIs in applications related to chronic diseases. First, we develop a prototype that focuses on AUIs in chronic disease-related applications. This prototype served as the foundation for our qualitative investigation into how individuals experience AUIs within the context of chronic diseases, through interviews and focus group studies. Additionally, all participants also completed a post-session survey to provide additional information on adaptations. During the same time, we conducted a quantitative survey to gather user preferences regarding various aspects of the adaptation process. By collecting and analysing qualitative and quantitative data and aligning with the existing literature, we established guidelines for designing AUIs for chronic disease-related applications. These guidelines were refined based on feedback obtained from a subsequent survey, resulting in the finalisation of nine guidelines.
We presented the initial findings of our user study at the 2024 International Conference on Software Engineering (ICSE) (Wang et al., 2024). In this paper, we significantly expand on our previous work by: (i) conducting a detailed and extended analysis of the emerging concepts identified in the user study, incorporating post-session survey data, and mapping the analysis between categories obtained from the study; (ii) enhancing our survey findings by including all collected data and conducting further analysis and interpretation; (iii) developing a set of new guidelines for designing AUIs in four stages; and (iv) refining our guidelines through an additional round of feedback collection. This work offers five key research contributions:
An investigation of the needs and preferences of users with chronic diseases for their mHealth app AUIs, expanding the adaptation literature by uncovering challenges in AUI usage and suggesting practical strategies for improved AUI design;
@@
Advancement of the discussion on this topic by presenting design trade-offs in AUIs when designing technology for users with chronic disease;
@@
Deeper understanding of how user preferences for different aspects of adaptations are influenced by diiferent demographic factors, including cultural backgrounds, contextual circumstances, Health condition and age.
@@
Development of comprehensive actionable guidelines for researchers, practitioners, and designers for designing AUIs in chronic disease domain; and
Initial evaluation of these guidelines with a cohort of 20 users and an iterative refinement of the guidelines.
@@
The rest of the paper is organised as follows. Section 2 provides a summary of our study’s motivation and summarises key related work. In Section 3, we present our research methodology. The results of qualitative and quantitative data analysis are presented in Sections 4 and 5, respectively. Section 6 offers an in-depth description of the proposed guidelines. We present the results from the feedback survey for the guidelines in Section 6.2. Finally, section 7 presents the limitations and threats to the validity of the research and Section 8 concludes this paper.
@@ 2. Motivation and Related Work
Adaptive User Interfaces: Due to the diversity of users and usage contexts, a User Interface (UI) developed for a fixed context of use may not be sufficient. AUIs adapt dynamically to users’ contexts and preferences, showing promise in addressing such issues. McTear (1993) defines an AUI as “a software artefact that improves its ability to interact with a user by constructing a user model based on partial experience with that user ”. Recent research highlights the importance of adaptable mHealth technologies in facilitating chronic disease self-management (McLean et al., 2011; Grua et al., 2020). An example are mobile applications developed for the management of diabetic hypoglycemia, which adapt the navigation system for hypoglycemic events (Pagiatakis et al., 2020). Under normal circumstances, the app displays a standard homepage for regular use (see Figure 1(a)). However, as depicted in Figure 1(b), access to other sections of the app will be restricted during a hypoglycemic event, and a prominent feature allows users to quickly contact emergency services. Jabeen et al. (2019) design an adaptable interface called PD-Helper to cater to the physical needs of Parkinson patients (see Figure 1(c)). The web page control section includes web elements that allow users to adjust font size, reload the page, open a new page, and navigate back to the main menu with a single tap on the screen.
@@
Chronic Disease Self-Management: Chronic illnesses pose a considerable health issue on a worldwide scale, accounting for a significant number of yearly deaths (WHO, 2022). These diseases are responsible for the deaths of 41 million individuals annually, representing 74% of all global fatalities (WHO, 2023). The prevalence of chronic diseases is steadily increasing, driven by the reclassification of previously fatal diseases as chronic diseases and the ageing of the population (Beaglehole et al., 2008). Despite advancements in medical science, significant challenges persist in finding cures for chronic diseases such as diabetes and Alzheimer’s disease (Hird et al., 2016). Traditional treatment paradigms do not address the multifaceted nature of chronic diseases (WHO, 2022; Hird et al., 2016), as treatment cannot be based solely on biological parameters, which require active engagement and self-management by patients (Deiss et al., 2006). Self-management entails actively engaging in self-care activities to enhance one’s behaviour and well-being (Deiss et al., 2006). Research has highlighted the efficacy of mHealth technology in supporting self-management of chronic diseases (Paré et al., 2007; Hamine et al., 2015). Although there are possible advantages of mHealth interventions for individuals with long-term health problems, studies indicate a trend where those who could benefit the most from these interventions tend to have lower utilisation rates (Han et al., 2010). A significant proportion of fatalities resulting from chronic diseases, accounting for 77%, occur in low and middle-income nations (WHO, 2023).
AUIs and Healthcare: Previous research has explored the application of AUIs in systems adapted for healthcare professionals (Eslami et al., 2018; Greenwood et al., 2003; Vogt and Meier, 2010). Eslami et al. (2018) conducted interviews and observations to investigate user preferences regarding data entry, language and vocabulary, information presentation, as well as providing help, warning, and feedback, with a primary focus on healthcare professionals. Similarly, Vogt and Meier (2010) examined AUI design issues that aim to simplify input and reduce the potential for errors, particularly in contexts such as smart hospitals. Greenwood et al. (2003) introduced a novel approach using reactive agents for AUIs in diabetes treatment decision support, customising data display according to clinician preferences. Despite the initial emphasis on healthcare professionals, there is an increasing recognition of the diverse user base of mHealth applications. Consequently, there has been an increasing number of literature describing various approaches, applications, and tools specifically tailored for patient-focused AUIs. Existing studies on AUI frameworks often focus on particular adaptive components or specific aspects of patient management (Shakshuki et al., 2015; Fröhlich et al., 2009; Yuan and Herbert, 2012). Shakshuki et al. (2015) proposed an AUI architecture for patient monitoring with a focus on health-related information adaptation. Fröhlich et al. (2009) explored the application of LoCa, primarily to monitor physiological data and the activity status of patients within a digital home environment, facilitating context-aware adaptation of workflows. Yuan and Herbert (2012) designed a fuzzy-logic-based context model for personalised healthcare services in chronic illness, prioritising health problem prediction and preventive measures based on user data rather than UI adaptation to individual user needs and context.
Despite the initial emphasis on healthcare professionals, there is an increasing recognition of the diverse user base of mHealth applications. AUIs have been deployed across a spectrum of mHealth applications, ranging from stroke rehabilitation (Burke et al., 2009), diabetes (Pagiatakis et al., 2020), cardiac disease (Mohan et al., 2008), dementia (Awada et al., 2018; Gullà et al., 2019) and Parkinson’s disease (Jabeen et al., 2019). These applications exhibit varying degrees of adaptability, ranging from the adjustment of difficulty levels of exercise activity (Burke et al., 2009) to the customisation of health-related information (Awada et al., 2018; Gullà et al., 2019), navigation adaptations (Pagiatakis et al., 2020; Awada et al., 2018), multimodal interfaces (Pagiatakis et al., 2020), information architecture (Mohan et al., 2008), and graphic design (Gullà et al., 2019; Mohan et al., 2008; Pagiatakis et al., 2020; Jabeen et al., 2019). However, most of these studies lack detailed explanations of their AUI development process, especially concerning the initial stages involving the collection of diverse end user requirements. Furthermore, the typical evaluation approach for AUIs focuses on overall application effectiveness without proper comparisons to non-adaptive UIs, which complicates drawing specific conclusions about the impact of AUIs. Although research in other domains suggests that AUIs can improve user performance and satisfaction compared to non-adaptive baselines (Gajos et al., 2006; Paymans et al., 2004; Tsandilas and Schraefel, 2005; Li and Gajos, 2014), disruptive adaptations, which alter users’ accustomed interaction patterns or break conventions, can result in frustration or dissatisfaction (Findlater and McGrenere, 2004; Rivera, 2005). Despite these insights, our understanding of how individuals with chronic diseases use AUIs and how to design mHealth technologies integrating AUIs to maximise benefits while minimising costs for this population remains limited.
@@
AUIs and Chronic Disease Self-Management: To enhance the successful deployment of mHealth applications, particularly for patients managing chronic conditions, it is imperative to tailor these applications to meet the needs of diverse user groups. There are well-recognised challenges to doing this, including: Firstly, chronic diseases have high heterogeneity, impacting patients differently in terms of triggers, symptoms, and severity (Harvey et al., 2012; Deiss et al., 2006; Audulv, 2013). Therefore, patients may have various self-management needs. Secondly, the design of UI must accommodate the changing phases of chronic diseases over time (di San Marco et al., 2019; Audulv, 2013). Many chronic conditions either worsen or improve with time, necessitating corresponding adjustments to self-management plans (Lorig and Holman, 2003). Furthermore, chronic diseases often co-morbid with other medical or psychological disorders, resulting in a broader spectrum of user characteristics and functionality requirements (di San Marco et al., 2019; Islam et al., 2014). For example, diabetes can lead to various complications such as vision loss, amputation, neuropathy, end-stage renal disease, cardiovascular disease, infections, and cognitive impairment (Vijan, 2015). Thirdly, the vast majority of chronic diseases are long-lasting and generally lifelong (WHO, 2022; Harvey et al., 2012), typically manageable but not curable. Therefore, mHealth technologies must sustain user engagement and motivation over the long term. In addition to the diverse nature of chronic diseases, patients also have a wide range of backgrounds, expertise, and demographic, psychological, and cognitive characteristics (Vasilyeva et al., 2005). While adapting interfaces can improve user acceptance and motivation (Duplaga, 2015; Han et al., 2010), existing AUI solutions often lack full adaptability, relying on predefined rules and overlooking unique characteristics of users (Grua et al., 2020; McLean et al., 2011). We wanted to explore the perspectives of users on AUIs in chronic disease management, seeking to establish guidelines for designing more effective AUIs that meet the varied needs of individuals with chronic diseases.

### v2

# stebox1linecolor=steBoxLine,linewidth=2pt,leftmargin=0cm,rightmargin=0cm,roundcorner=5pt, topline=false,bottomline=false,rightline=false,leftline=true,backgroundcolor=gray!10 Designing Adaptive User Interfaces for mHealth Applications Targeting Chronic Disease: A User-Centered Approach

@@ Abstract.
Mobile Health (mHealth) applications have demonstrated considerable potential in supporting chronic disease self-management; however, they remain underutilized due to low engagement, limited accessibility, and poor long-term adherence. These issues are particularly prominent among users with chronic disease, whose needs and capabilities vary widely. To address this, Adaptive User Interfaces (AUIs) offer a dynamic solution by tailoring interface features to users’ preferences, health status, and contexts. This paper presents a two-stage study to develop and validate actionable AUI design guidelines for mHealth applications. In stage one, an AUI prototype was evaluated through focus groups, interviews, and a standalone survey, revealing key user challenges and preferences. These insights informed the creation of an initial set of guidelines. In stage two, the guidelines were refined based on feedback from 20 end users and evaluated by 43 software practitioners through two surveys. This process resulted in nine finalized guidelines. To assess real-world relevance, a case study of four mHealth applications was conducted, with findings supported by user reviews highlighting the utility of the guidelines in identifying critical adaptation issues. This study offers actionable, evidence-based guidelines that help software practitioners design AUI in mHealth to better support individuals managing chronic diseases.
@@ 1. Introduction
Chronic diseases, including conditions such as asthma, heart disease, and diabetes, present formidable challenges to healthcare systems around the world (WHO, 2022). The management of these long-term health conditions transcends simple medical treatment, with an increasing emphasis on empowering patients to actively engage in self-management practices (WHO, 2022). The use of Mobile Health (mHealth) applications has emerged as a promising avenue for promoting self-management by strengthening medication adherence and facilitating self-tracking capabilities (Hamine et al., 2015). Despite their potential, studies indicate that many individuals who could benefit from mHealth technology do not fully engage with these tools, particularly in developing countries where non-adherence rates tend to be higher (Han et al., 2010; Beaglehole et al., 2008). Increasing mHealth adoption among chronic disease patients requires greater customization and flexibility (Choe et al., 2017; Han et al., 2010), yet several design and implementation challenges must first be addressed. Due to their inherent heterogeneity, chronic diseases affect individuals in various ways and can co-occur with other medical or psychological disorders, making self-management more complex (Harvey et al., 2012; Audulv, 2013; di San Marco et al., 2019; Audulv, 2013). In addition, chronic diseases typically persist throughout a person’s lifetime (WHO, 2022; Harvey et al., 2012), mHealth applications need to be designed to sustain user engagement and motivation over the long term.
One critical area of research that addresses barriers to sustained use of technology is accessibility. Traditionally, accessibility research has concentrated on meeting the needs of individuals with disabilities such as blindness, low vision, and physical impairment, dominating over half of the studies conducted in the last decade (Zhou et al., 2020; Mack et al., 2022, 2021; Depuy et al., 2023). Chronic diseases, which affect a significant portion of the population, remain underrepresented in accessibility research, highlighting the need for solutions to address the diverse challenges faced by individuals managing these conditions. Adaptive User Interfaces (AUIs) offer a promising solution to bridge this gap by dynamically tailoring the User Interface (UI) to accommodate the unique needs, goals, and contextual circumstances of each individual (Norcio and Stanley, 1989; Firmenich et al., 2019). In this work, we focus on adaptation, which broadly refers to interface modifications, encompassing both system-driven changes (adaptivity) and user-controlled customizations (adaptability) (Norcio and Stanley, 1989). Additionally, we acknowledge mixed-initiative adaptation, where both the user and the system collaboratively share the responsibility of adapting the interface (Horvitz, 1999; Mezhoudi and Vanderdonckt, 2021). Despite increasing interest in employing AUIs in chronic disease-related applications (Setiawan et al., 2019; Talboom-Kamp et al., 2018), they often overlook various user characteristics and interactions (Grua et al., 2020; McLean et al., 2011). Most chronic disease-related applications view design adaptations as a complex task, requiring expertise in psychology, physiology, human behavior, user experience analysis and interpretation with regard to the underlying behavior and health status (Grua et al., 2020; Vasilyeva et al., 2005). Many software practitioners lack the expertise required to implement theories and models, making it challenging to access the necessary skills for mHealth-related projects (Chib and Lin, 2018). Moreover, there is a lack of established resources and guidelines for AUI development (Zimmerman et al., 2020). Existing studies on the development of AUI within these applications reveal a substantial deficiency in the literature concerning the foundational stages of AUI design (Wang et al., 2023b). This gap consequently limits the understanding of how AUIs are perceived and utilized by individuals with chronic diseases, potentially resulting in the underutilization of the benefits that adaptive systems can offer (Kornfield et al., 2020). Bridging this gap is therefore essential for optimizing AUI design and improving user engagement in chronic disease-related applications.
In this work, we present a two-stage approach aimed at designing and validating mHealth adaptation guidelines for chronic disease-related applications. In stage one, we developed an AUI prototype and collected input from chronic disease patients through focus groups, interviews, and surveys, forming the initial design guidelines. In stage two, we refined these guidelines based on end-user and software practitioner feedback, and subsequently validated through real-world mHealth applications. We presented the initial findings of user study in stage one at the 2024 International Conference on Software Engineering (ICSE) (Wang et al., 2024). In this work, we build on and significantly extend our earlier conference paper by advancing from stage one to stage two. Specifically, we: 1) enhance our user survey findings in stage one by including all collected data and conducting further analysis and interpretation; 2) develop a set of new guidelines for designing AUIs for mHealth application targeting chronic diseases; 3) refine these guidelines through an additional round of feedback from end-users and software practitioners; and 4) conduct a case study on existing mHealth applications to further validate the guidelines. This work offers five key research contributions:
@@
Advancement of the discussion on this topic by presenting design trade-offs in AUI when designing technology for users with chronic disease;
@@
Deeper understanding of how user preferences for different aspects of adaptations are influenced by different demographic factors, including cultural backgrounds, contextual circumstances, Health condition and age;
@@
Development of comprehensive actionable guidelines for researchers, practitioners, and designers for designing AUI in the chronic disease domain;
Multi-stage evaluation and refinement of actionable guidelines through feedback collection from end-users and software practitioners;
Validation of refined guidelines through a case study on real-world mHealth applications, demonstrating their practical applicability and effectiveness in addressing adaptation challenges in the chronic disease domain.
@@
The rest of the paper is organized as follows. Section 2 provides a summary of our study’s motivation and summarizes key related work. In Section 3, we present our research methodology. The results of the qualitative and quantitative analysis of stage one are detailed in Sections 4 and 5, respectively. In stage two, Section 6.2 presents the evaluation results of the guidelines from an end-user feedback survey. Section 6.3 discusses the feedback from software practitioners on the refined guidelines. Furthermore, Section 6.5 evaluates the practical applicability of the guidelines through a case study involving real-world mHealth applications. The finalized set of guidelines is outlined in Section 6.4. Lastly, Section 7 discusses threats to validity of the study and Section 8 concludes the paper.
@@ 2. Motivation and Related Work
2.1. Chronic Disease Self-Management
@@
Responsible for 41 million deaths annually and representing 74% of all global deaths (WHO, 2022, 2023), chronic disease poses a major global health challenge. The prevalence of chronic diseases is steadily increasing, driven by the reclassification of previously fatal diseases as chronic diseases and the aging of the population (Beaglehole et al., 2008). Traditional treatment paradigms do not address the multifaceted nature of chronic diseases (WHO, 2022; Hird et al., 2016), as treatment cannot be based solely on biological parameters, which require active engagement and self-management by patients (Deiss et al., 2006). Self-management involves actively participating in self-care activities to improve behavior and well-being (Deiss et al., 2006). Research has highlighted the efficacy of mHealth applications in supporting the self-management of chronic diseases (Paré et al., 2007; Hamine et al., 2015). However, despite the potential benefits of such interventions, evidence suggests that those who stand to gain the most often exhibit lower levels of engagement and adoption (Han et al., 2010). This disparity is especially concerning given that approximately 77% of chronic disease-related fatalities occur in low- and middle-income countries, where access to consistent and effective healthcare remains limited (WHO, 2023).
@@
There are well-recognized challenges in designing mHealth technology for chronic disease self-management, one of which is the high heterogeneity of chronic diseases. These diseases affect patients differently in terms of triggers, symptoms, and severity (Harvey et al., 2012; Deiss et al., 2006; Audulv, 2013), resulting in a wide range of self-management needs across individuals. Secondly, the design of the UI must account for the evolving nature of chronic diseases over time (di San Marco et al., 2019; Audulv, 2013). As many chronic diseases fluctuate, the corresponding adjustments to self-management strategies are required (Lorig and Holman, 2003). Furthermore, chronic diseases are often co-morbid with other medical or psychological disorders, resulting in a broader spectrum of user characteristics and functionality requirements (di San Marco et al., 2019; Islam et al., 2014). For example, diabetes can cause various complications, such as vision loss, amputation, neuropathy, end-stage renal disease, cardiovascular disease, infections, and cognitive impairment (Vijan, 2015). Thirdly, the vast majority of chronic diseases are long-lasting and generally lifelong (WHO, 2022; Harvey et al., 2012), generally manageable but not curable. Therefore, mHealth applications must sustain user engagement and motivation over the long term. In addition to the diverse nature of chronic diseases, patients also have a wide range of backgrounds, expertise, and demographic, psychological, and cognitive characteristics (Vasilyeva et al., 2005). While adapting interfaces can improve user acceptance and motivation (Duplaga, 2015; Han et al., 2010), existing AUI solutions often have limited adaptation options, relying on predefined rules and overlooking unique characteristics of users (Grua et al., 2020; McLean et al., 2011). Creating mHealth applications that offer access to knowledge and information is essential to prevent physical and social disparities (Han et al., 2010), especially in developing countries where nonadherence to treatment remains a significant issue (Beaglehole et al., 2008).
2.1.1. Chronic disease-related applications and accessibility study
While multidisciplinary efforts have contributed to chronic disease-related applications, many technological solutions remain anchored in medicalized perspectives, often viewing users solely as patients rather than individuals with various priorities and lifestyles (Burgess et al., 2019; Huh and Ackerman, 2012). A considerable amount of research exists on the effect of these mHealth applications on treatment regimen adherence (e.g. (Hamine et al., 2015; Gandapur et al., 2016)), application design features (e.g. (Sockolow et al., 2021; Matthew-Maich et al., 2016)), and the evaluation of mHealth applications (e.g. (Yang et al., 2020)). Some exceptions exist in exploring design strategies for individuals with chronic diseases (e.g. (Mack et al., 2022; Paymal and Homewood, 2024; Epstein et al., 2016)). These initiatives aim to explore the evolving nature of chronic diseases and how changes in users’ physical, cognitive, and emotional needs affect their interaction with technology. However, many self-management tools continue to generalize the patient experience and overlook the complexity of managing health in everyday contexts (Nunes et al., 2015). To address this gap, researchers have called for greater customization and personalization in mHealth applications design, advocating systems that can adapt to users’ unique health trajectories and life contexts (Nunes et al., 2015; Ryu et al., 2023; McLean et al., 2011; Grua et al., 2020). A critical domain that aligns with these goals is accessibility research, which traditionally focuses on improving system usability for individuals with disabilities (Radcliffe et al., 2021; Zhou et al., 2020; Alshayban et al., 2020; Daihua et al., 2015), including those with visual (Depuy et al., 2023; Alshayban et al., 2020; Milne et al., 2014) and cognitive impairments (Yu et al., 2017), or those of low socioeconomic backgrounds (Stowell et al., 2018). Chronic diseases remain underrepresented in accessibility studies, despite the growing evidence of substantial accessibility challenges in self-management applications (Kim, 2019; Mack et al., 2021). Researchers argue that individuals with chronic diseases encounter nuanced challenges not fully captured by the general disability frameworks (Mack et al., 2022). In this context, AUI offers significant potential in mHealth applications by enhancing both accessibility and usability. By tailoring interfaces to the diverse capabilities and preferences of users, AUI ensures that these applications are not only functional but also inclusive (Wang et al., 2023b).
2.2. Adaptive User Interfaces
Given the diversity of users and contexts, fixed-context UIs are often inadequate; AUIs, which dynamically adjust to user preferences and environments, offer a promising solution. McTear (1993) defines an AUI as “a software artefact that improves its ability to interact with a user by constructing a user model based on partial experience with that user”. Recent research highlights the importance of adaptive mHealth applications in facilitating chronic disease self-management (Nunes et al., 2015; Ryu et al., 2023; McLean et al., 2011; Grua et al., 2020). For example, in the treatment of diabetic hypoglycemia, an application developed by Pagiatakis et al. (2020) adapts its navigation system during hypoglycemic events. Under normal conditions, the application displays a standard homepage for everyday use (see Figure 1(a)). However, during a hypoglycemic episode, as shown in Figure 1(b), the application restricts access to non-essential sections and prominently features a quick access emergency contact button to ensure user safety. Similarly, Jabeen et al. (2019) designed PD-Helper, AUIs tailored to support individuals with Parkinson’s disease (see Figure 1(c)). The application provides customization control, allowing users to adjust font size, refresh pages, open new tabs, and return to the main menu with a single tap, thus addressing the motor limitations commonly experienced by these patients.
2.2.1. AUIs and healthcare
Previous research has explored the application of AUIs in systems adapted for healthcare professionals (Eslami et al., 2018; Greenwood et al., 2003; Vogt and Meier, 2010). Eslami et al. (2018) conducted interviews and observations to investigate user preferences regarding data entry, language and vocabulary, and information presentation, as well as providing help, warning, and feedback, with a primary focus on healthcare professionals. Similarly, Vogt and Meier (2010) examined AUI design issues that aim to simplify input and reduce the potential for errors, particularly in contexts such as smart hospitals. Greenwood et al. (2003) introduced a novel approach using reactive agents for AUIs in diabetes treatment decision support, customizing data display according to clinician preferences. Despite the initial emphasis on healthcare professionals, there is an increasing recognition of the diverse user base of mHealth applications. Consequently, there has been an increasing number of studies describing various approaches, applications, and tools specifically tailored for patient-focused AUIs. Existing studies on AUI frameworks often focus on particular adaptive components or specific aspects of patient management (Shakshuki et al., 2015; Fröhlich et al., 2009; Yuan and Herbert, 2012). Shakshuki et al. (2015) proposed an AUI architecture for patient monitoring with a focus on health-related information adaptation. Fröhlich et al. (2009) explored the application of the LoCa (A Location and Context-aware eHealth Infrastructure) project, primarily to monitor physiological data and the activity status of patients within a digital home environment, facilitating context-aware adaptation of workflows. Yuan and Herbert (2012) designed a fuzzy-logic-based context model for personalized healthcare services in chronic diseases, prioritizing the prediction of health problems and preventive measures based on user data rather than UI adaptation to individual user needs and context.
AUIs have been deployed across a spectrum of mHealth applications, ranging from stroke rehabilitation (Burke et al., 2009), diabetes (Pagiatakis et al., 2020), cardiac disease (Mohan et al., 2008), dementia (Awada et al., 2018; Gullà et al., 2019), and Parkinson’s disease (Jabeen et al., 2019). These applications exhibit different adaptations, ranging from the adjustment of the difficulty levels of exercise activity (Burke et al., 2009) to the customization of health-related information (Awada et al., 2018; Gullà et al., 2019), navigation adaptations (Pagiatakis et al., 2020; Awada et al., 2018), multimodal interfaces (Pagiatakis et al., 2020), information architecture (Mohan et al., 2008) and graphic design (Gullà et al., 2019; Mohan et al., 2008; Pagiatakis et al., 2020; Jabeen et al., 2019). Based on the findings of the previous Systematic Literature Review (SLR) by Wang et al. (2023b), existing user models, which represent various user dimensions to support adaptation, predominantly utilize physical and physiological characteristics to generate AUI. Most existing studies acquire user data through methods such as user questionnaires or by allowing users to manually adjust settings and preferences during application usage. Regarding the adaptation mechanism, studies predominantly use rule-based adaptation and predictive algorithm-based adaptation to adapt interfaces to user needs. However, most of these studies lack detailed explanations of their AUI development process, especially with regard to the initial stages that involve the collection of diverse end-user requirements. Furthermore, the typical evaluation approach for AUIs focuses on overall application effectiveness without proper comparisons to non-adaptive UIs, which complicates drawing specific conclusions about the impact of AUIs. Although research in other domains suggests that AUIs can improve user performance and satisfaction compared to non-adaptive baseline (Gajos et al., 2006; Paymans et al., 2004; Tsandilas and Schraefel, 2005; Li and Gajos, 2014), disruptive adaptations, which alter user accustomed interaction patterns or break conventions, can result in frustration or dissatisfaction (Findlater and McGrenere, 2004; Rivera, 2005). Despite these insights, our understanding of how individuals with chronic diseases use AUIs and how to design mHealth applications that integrate AUIs to maximize benefits while minimizing costs for this population remains limited.